The shaft of a steam turbine produces 600 Nm torque at 50rev/s. Calculate the work transfer rate from the steam


1
Expert's answer
2022-04-14T08:55:03-0400
"P=\tau\times \omega"

"P=600\times 50\times 6.28=1.88*10^5\:\rm W"